OpenStudy (anonymous):

Need help. Find the functions of f and g and find the domain. a. (f+g) (x) B. (f-g) (x) C.( f*g)(x) D. (F/g) (x) Problem F(x) = 3x +4 g(x)=2x-3

OpenStudy (anonymous):

(f+g)(x) = f(x) + g(x) = 5x + 1, and so on

OpenStudy (anonymous):

Where did the 4 go? I got 5x +1x+4. All I have to do to solve is add both equations to gether?